# Turbocharging constraints on dark matter substructure through a synthesis of strong lensing flux ratios and extended lensed arcs

### v2

@@ Abstract
Strong gravitational lensing provides a purely gravitational means to infer properties of dark matter halos and thereby constrain the particle nature of dark matter. Strong lenses sometimes appear as four lensed images of a background quasar accompanied by spatially-resolved emission from the quasar host galaxy encircling the main deflector (lensed arcs). We present methodology to simultaneously reconstruct lensed arcs and relative image magnifications (flux ratios) in the presence of full populations of dark matter subhalos and line-of-sight halos. To this end, we develop a new approach for multi-plane ray tracing that accelerates lens mass and source light reconstruction by factors of $\sim 100-1000$. Using simulated data with a cold dark matter (CDM) ground truth, we show that simultaneous reconstruction of lensed arcs and flux ratios isolates small-scale perturbations to flux ratios by dark matter substructure from uncertainties associated with the main deflector mass profile on larger angular scales. Relative to analyses that use only image positions and flux ratios to constrain the lens model, incorporating arcs strengthens likelihood ratios penalizing warm dark matter (WDM) with a characteristic suppression scale $m_{\rm{hm}}/M_{\odot}$ in the range $\left[10^{7}-10^{7.5}\right]$, $\left[10^{7.5}-10^{8}\right]$, $\left[10^{8}-10^{8.5}\right]$, $\left[10^{8.5}-10^{9}\right]$ by factors of $1.3$, $2.5$, $5.6$, and $13.1$, respectively, and the $95\%$ exclusion limit improves by 0.5 dex in $\log_{10}m_{\rm{hm}}$. The enhanced sensitivity to low-mass halos enabled by these methods pushes the observational frontier of substructure lensing to the threshold of galaxy formation, enabling stringent tests of any theory that alters the properties of dark matter halos.
@@
pubyear: 2024pagerange: Turbocharging constraints on dark matter substructure through a synthesis of strong lensing flux ratios and extended lensed arcsA
@@ 1 Introduction
The abundance and internal structure of dark matter halos depends on the particle nature of dark matter (Buckley & Peter, 2018). As such, characterizing the properties of dark matter substructure, the low-mass ($<10^{10}M_{\odot}$) halos that surround galaxies and permeate the cosmos, enables tests of fundamental dark matter physics. On sub-galactic scales, differences between the concordance cosmological model of cold dark matter (CDM) diverge from the predictions of other theories. For example, if the dark matter has a sufficiently large ($\gtrsim 1\rm{kpc}$) free-streaming length, the abundance and central density of dark matter halos become suppressed on scales comparable to that of a low-mass galaxy, relative to CDM (Bond & Szalay, 1983; Bode et al., 2001; Bose et al., 2016; Lovell, 2020; Stücker et al., 2022). In this class of theory, categorically referred to as warm dark matter (WDM), the free-streaming scale depends on the formation mechanism and mass of the dark matter particle (e.g. Schneider et al., 2012; Abazajian & Kusenko, 2019). Alternatively, theories with self-interacting dark matter (SIDM) posit that the dark matter behave as CDM on cosmological scales, but experiences self-interactions inside high-density regions, such as dark matter halos (Spergel & Steinhardt, 2000; Tulin et al., 2013). In SIDM, halos initially form a central core, and eventually undergo a process referred to as core collapse that causes an order-of-magnitude increase in their central density of halos (Balberg et al., 2002; Gilman et al., 2021; Gilman et al., 2023; Nadler et al., 2023; Yang et al., 2023).
@@
In search of new physics, observational probes of dark matter structure from dwarf galaxies (Kim et al., 2018; Correa, 2021; Nadler et al., 2021; Bechtol et al., 2022; Dekker et al., 2022; Akita & Ando, 2023; Slone et al., 2023; Nadler et al., 2024), stellar streams (Bovy et al., 2017; Banik et al., 2018; Bonaca et al., 2019; Banik et al., 2021), and gravitational lensing (Dalal & Kochanek, 2002; Vegetti et al., 2014; Nierenberg et al., 2014; Hezaveh et al., 2016; Birrer et al., 2017; Nierenberg et al., 2017; Vegetti et al., 2018; Gilman et al., 2020, 2021; He et al., 2022; Sengül et al., 2022; Wagner-Carena et al., 2023; Dike et al., 2023; Powell et al., 2023; Dhanasingham et al., 2023; Keeley et al., 2023; Gilman et al., 2023; Mondino et al., 2023; Nightingale et al., 2024) characterize the properties of substructure on sub-galactic scales. If the dark matter has no coupling to the standard model besides gravity, these cosmic probes of dark matter constitute the only experiments with which to investigate its properties (see the reviews by Drlica-Wagner et al., 2022; Bechtol et al., 2022).
@@
Among the various observational probes that constrain dark matter properties through studies of low-mass halos, strong gravitational lensing provides the unique capability to characterize the properties of substructure across cosmological distances, and across several Gyr of cosmic time (see the recent review by Vegetti et al. (2023)). Strong lensing refers to a phenomenon in which multiplehighly-magnified and distorted images of a background source appear due to deflection of light around an intervening cosmic structure,such as a galaxy. As lensing depends only on gravity, it circumvents systematic uncertainties associated with using luminous matter as a tracer for the underlying dark matter, and can characterize both the abundance and internal structure of dark matter halos (Minor et al., 2021; Amorisco et al., 2022; Gilman et al., 2022; Ballard et al., 2023). Through the direct, purely gravitational detection of dark halos, strong lensing can extend the reach of cosmic probes of dark matter structure to scales below the threshold of galaxy formation $\sim 10^{7}M_{\odot}$.
@@
Performing such a measurement requires exquisite data. Over the last decade, radio interferometry (Koopmans et al., 2004; McKean et al., 2007; Spingola et al., 2018), the Hubble Space Telescope (HST) (Shajib et al., 2019; Nierenberg et al., 2017, 2020), the W.M. Keck Observatory (Nierenberg et al., 2014), and the James Webb Space Telescope (JWST) (Nierenberg et al., 2023) have observed a particular class of strong lens system in which a quasar becomes quadruply-imaged. These systems are ideally suited to probe low-mass dark matter structure because the relative magnifications among lensed images (flux ratios) experience strong perturbation by low-mass halos. The minimum mass sensitivity of the flux ratios is determined by the size of the lensed background source (Dobler & Keeton, 2006), and recent JWST observations of the “warm dust region” around the background quasar are expected to provide sensitivity to halos at $\lesssim 10^{7}M_{\odot}$.
@@
As shown in Figure 1, a quadruply-imaged quasar can sometimes appear alongside spatially-resolved lensed emission from the quasar host galaxy, or lensed arcs. While the flux ratios provide sensitive localized constraints on the lens model, the lensed arcs that encircle the main deflector impose stringent constraints on the mass profile across larger angular scales (e.g. Shajib et al., 2020; Powell et al., 2022). Incorporating constraints from the arcs leads to tighter constraints on the main deflector mass profile, improving the precision of model-predicted flux ratios (Oh et al., in prep). However, due mainly to computational limitations, no existing methodology enables the self-consistent reconstruction of lensed arcs and quasar flux ratios in the presence of potentially tens-of-thousands of dark matter halos. As a result, substructure lensing analyses performed with quadruple-image lenses use only the image positions and flux ratios to constrain the lens model and the properties of dark matter substructure.
@@ 2.1 The Bayesian inference problem
where $\boldsymbol{O}=\left(\boldsymbol{O_{1}},\boldsymbol{O_{2}},....\right)$ represents the data from a sample of lenses, $\pi\left({{\bf{q}}}\right)$ represents the prior on ${{\bf{q}}}$, and where $\mathcal{L}\left({{\bf{O}}}_{i}|{{\bf{q}}}\right)$ represents the likelihood function for a single lens with data ${{\bf{O}}}_{i}\equiv\left(\boldsymbol{O_{\rm{pos}}},\boldsymbol{O_{\rm{img}}},\boldsymbol{O_{\rm{f}}}\right)$. To connect the hyper-parameters ${{\bf{q}}}$ with $\boldsymbol{O_{i}}$, we must generate realizations, ${\boldsymbol{m}_{\rm{sub}}}$, of dark matter substructure from the model specified by ${{\bf{q}}}$. Hereafter, a realization of substructure refers to the set of halo masses, positions, and density profiles from which we can compute a lensing deflection field. The likelihood follows from marginalizing over many realizations
@@
In the preceding equation, we have introduced a vector of nuisance parameters, $\boldsymbol{\mathcal{N}}$, which include quantities that parameterize the main deflector mass profile (hereafter the macromodel), the spatial extent and structure of the lensed background quasar, and the surface brightness profile of the quasar host galaxy. The likelihood appearing in the integrand is the product of an astrometric likelihood computed with the lensed image positions $\mathcal{L}\left(\boldsymbol{O}_{\rm{pos}}|{\boldsymbol{m}_{\rm{sub}}},\boldsymbol{\mathcal{N}}\right)$, the imaging data likelihood $\mathcal{L}\left(\boldsymbol{O}_{\rm{img}}|{\boldsymbol{m}_{\rm{sub}}},\boldsymbol{\mathcal{N}}\right)$, and the flux ratio likelihood $\mathcal{L}\left(\boldsymbol{O}_{\rm{fr}}|{\boldsymbol{m}_{\rm{sub}}},\boldsymbol{\mathcal{N}}\right)$.
@@ 2.2.2 The flux ratio likelihood
In our analysis, we generate $\sim 10^{6}$ realizations per lens, and accept the top $N_{\rm{accept}}=3,000$ samples corresponding to the lowest values of $S$. This results in values of $\epsilon$ that range between 0.01 and 0.1 among the (mock) lenses in our sample. As in previous work, we handle observational measurement uncertainties by adding them post-processing to the model-predicted flux ratios before computing $S$. The choices we implement regarding the $\epsilon$ tolerance we use is motivated by the convergence tests conducted for ABC rejection algorithms applied to image flux ratios by Gilman et al. (2020). Finally, we obtain a continuous approximation of the flux-ratio likelihood function, $\mathcal{L}\left(\boldsymbol{O_{\rm{f}}}|{\boldsymbol{m}_{\rm{sub}}},\boldsymbol{\mathcal{N}}\right)$, by applying a Gaussian kernel density estimator to the accepted samples.
@@ 2.2.3 The imaging data likelihood
We compute the imaging data likelihood by applying an ABC rejection cut to a second summary statistic, $L$, which we define as the likelihood of the imaging data $L\equiv\mathcal{L}\left(\boldsymbol{O_{\rm{img}}}|{\boldsymbol{m}_{\rm{sub}}},\boldsymbol{\mathcal{N}}\right)$. In principle, we could incorporate the imaging data likelihood directly through importance sampling with weights equal to $L$, but this causes numerical stability issues when the relative likelihood among the accepted samples fluctuate by large factors of $\sim e^{5}$ due to the high dimension of the data vector. We define the acceptance threshold on $L$ as the value of $L$ that corresponds to the top $2$nd percentile of the imaging data likelihoods. This choice represents a compromise between selecting models that best fit the imaging data, and accepting realizations that minimize the $S$ statistic computed with respect to the flux ratios.
As we show in Section 5, the imaging data enables stronger constraints on substructure properties by imposing tighter constraints on the mass profile of the main deflector than one obtains from analyzing only the image positions and flux ratios. However, our method of incorporating imaging data exposes the likelihood function to a systematic bias associated with reconstruction of the source light profile with substructure in the lens model. This bias manifests as a systematic preference for lens models with less substructure. Moreover, in rare cases a dark matter halo imparts a strong perturbation to the surface brightness of a lensed arc, a feature of the data that the inference methodology we present in this work is not intended to properly capture in the likelihood. For reasons we elaborate on further in Appendix A, we conclude that the imaging data likelihood will not give a reliable inference of substructure properties on its own when calculated according to the methodology we present in this work.
To mitigate the effects of systematics associated with the imaging data and source reconstruction in the presence of substructure, we demand that our posterior distribution have the property
@@
That is, the posterior should equal the prior on ${{\bf{q}}}$ when computed only using the imaging data (as $\epsilon\rightarrow\infty$ the flux ratio ratio likelihood term becomes uninformative). We can arrange that our posterior meets this requirement by introducing importance sampling weights, $w$, given by
@@
which is simply the inverse of the likelihood of measuring $\boldsymbol{\mathcal{O}_{\rm{img}}}$ for a given population of substructure ${\boldsymbol{m}_{\rm{sub}}}$ generated from the model $\boldsymbol{\rm{q}}$. This term effectively introduces a prior on the realizations ${\boldsymbol{m}_{\rm{sub}}}$ based on the source reconstruction. As one can easily verify, applying these importance weights and performing a dark matter inference without incorporating constraints from the flux ratios yields a posterior distribution equal to $\pi\left({{\bf{q}}}\right)$.

### v5

@@ Abstract
Strong gravitational lensing provides a purely gravitational means to infer properties of dark matter halos and thereby constrain the particle nature of dark matter. Strong lenses sometimes appear as four lensed images of a background quasar accompanied by spatially-resolved emission from the quasar host galaxy encircling the main deflector (lensed arcs). We present methodology to simultaneously reconstruct lensed arcs and relative image magnifications (flux ratios) in the presence of full populations of subhalos and line-of-sight halos. To this end, we develop a new approach for multi-plane ray tracing that accelerates lens mass and source light reconstruction by factors of $\sim 100-1000$. Using simulated data, we show that simultaneous reconstruction of lensed arcs and flux ratios isolates small-scale perturbations to flux ratios by dark matter substructure from uncertainties associated with the main deflector mass profile on larger angular scales. Relative to analyses that use only image positions and flux ratios to constrain the lens model, incorporating arcs strengthens likelihood ratios penalizing warm dark matter (WDM) with a suppression scale $m_{\rm{hm}}/M_{\odot}$ in the range $\left[10^{7}-10^{7.5}\right]$, $\left[10^{7.5}-10^{8}\right]$, $\left[10^{8}-10^{8.5}\right]$, $\left[10^{8.5}-10^{9}\right]$ by factors of $1.3$, $2.5$, $5.6$, and $13.1$, respectively, for a cold dark matter (CDM) ground truth. The $95\%$ exclusion limit improves by 0.5 dex in $\log_{10}m_{\rm{hm}}$. The enhanced sensitivity to low-mass halos enabled by these methods pushes the observational frontier of substructure lensing to the threshold of galaxy formation, enabling stringent tests of any theory that alters the properties of dark matter halos.
@@
pubyear: 2024pagerange: Turbocharging constraints on dark matter substructure through a synthesis of strong lensing flux ratios and extended lensed arcs23
@@ 1 Introduction
The abundance and internal structure of dark matter halos depends on the particle nature of dark matter (Buckley & Peter, 2018). As such, characterizing the properties of dark matter substructure, the low-mass ($<10^{10}M_{\odot}$) halos that surround galaxies and permeate the cosmos, enables tests of fundamental dark matter physics. On sub-galactic scales, differences between the concordance cosmological model of cold dark matter (CDM) diverge from the predictions of other theories. For example, if the dark matter has a sufficiently large ($\gtrsim 1\rm{kpc}$) free-streaming length, the abundance and central density of dark matter halos become suppressed on scales comparable to that of a low-mass galaxy, relative to CDM (Bond & Szalay, 1983; Bode et al., 2001; Bose et al., 2016; Lovell, 2020; Stücker et al., 2022). In this class of theory, categorically referred to as warm dark matter (WDM), the free-streaming scale depends on the formation mechanism and mass of the dark matter particle (e.g. Schneider et al., 2012; Abazajian & Kusenko, 2019). Alternatively, theories with self-interacting dark matter (SIDM) posit that the dark matter behave as CDM on cosmological scales, but experiences self-interactions inside high-density regions, such as dark matter halos (Spergel & Steinhardt, 2000; Tulin et al., 2013). In SIDM, halos initially form a central core, and eventually undergo a process referred to as core collapse that causes an order-of-magnitude increase in their central density of halos (Balberg et al., 2002; Gilman et al., 2021; Gilman et al., 2023; Nadler et al., 2023b; Yang et al., 2023).
@@
In search of new physics, observational probes of dark matter structure from dwarf galaxies (Kim et al., 2018; Correa, 2021; Nadler et al., 2021; Bechtol et al., 2022; Dekker et al., 2022; Akita & Ando, 2023; Slone et al., 2023; Nadler et al., 2024), stellar streams (Bovy et al., 2017; Banik et al., 2018; Bonaca et al., 2019; Banik et al., 2021), and gravitational lensing (Dalal & Kochanek, 2002; Vegetti et al., 2014; Nierenberg et al., 2014; Hezaveh et al., 2016; Birrer et al., 2017; Nierenberg et al., 2017; Vegetti et al., 2018; Gilman et al., 2020, 2021; He et al., 2022; Sengül et al., 2022; Wagner-Carena et al., 2023; Dike et al., 2023; Powell et al., 2023; Dhanasingham et al., 2023; Keeley et al., 2023; Gilman et al., 2023; Mondino et al., 2024; Nightingale et al., 2024) characterize the properties of substructure on sub-galactic scales. If the dark matter has no coupling to the standard model besides gravity, these cosmic probes of dark matter constitute the only experiments with which to investigate its properties (see the reviews by Drlica-Wagner et al., 2022; Bechtol et al., 2022).
@@
Among the various observational probes that constrain dark matter properties through studies of low-mass halos, strong gravitational lensing provides the unique capability to characterize the properties of substructure across cosmological distances, and across several Gyr of cosmic time (see the recent review by Vegetti et al. (2023)). Strong lensing refers to a phenomenon in which multiple highly-magnified and distorted images of a background source appear due to deflection of light around an intervening cosmic structure, such as a galaxy. As lensing depends only on gravity, it circumvents systematic uncertainties associated with using luminous matter as a tracer for the underlying dark matter, and can characterize both the abundance and internal structure of dark matter halos (Minor et al., 2021; Amorisco et al., 2022; Gilman et al., 2022; Ballard et al., 2024). Through the direct, purely gravitational detection of dark halos, strong lensing can extend the reach of cosmic probes of dark matter structure to scales below the threshold of galaxy formation $\sim 10^{7}M_{\odot}$.
@@
Performing such a measurement requires exquisite data. Over the last decade, radio interferometry (Koopmans et al., 2004; McKean et al., 2007; Spingola et al., 2018), the Hubble Space Telescope (HST) (Shajib et al., 2019; Nierenberg et al., 2017, 2020), the W.M. Keck Observatory (Nierenberg et al., 2014), and the James Webb Space Telescope (JWST) (Nierenberg et al., 2024) have observed a particular class of strong lens system in which a quasar becomes quadruply-imaged. These systems are ideally suited to probe low-mass dark matter structure because the relative magnifications among lensed images (flux ratios) experience strong perturbation by low-mass halos. The minimum mass sensitivity of the flux ratios is determined by the size of the lensed background source (Dobler & Keeton, 2006), and recent JWST observations of the “warm dust region” around the background quasar are expected to provide sensitivity to halos at $\lesssim 10^{7}M_{\odot}$.
@@
As shown in Figure 1, a quadruply-imaged quasar can sometimes appear alongside spatially-resolved lensed emission from the quasar host galaxy, or lensed arcs. While the flux ratios provide sensitive localized constraints on the lens model, the lensed arcs that encircle the main deflector impose stringent constraints on the mass profile across larger angular scales (e.g. Shajib et al., 2020; Powell et al., 2022). Incorporating constraints from the arcs leads to tighter constraints on the main deflector mass profile, improving the precision of model-predicted flux ratios (Oh et al., 2024). However, due mainly to computational limitations, no existing methodology enables the self-consistent reconstruction of lensed arcs and quasar flux ratios in the presence of potentially tens-of-thousands of dark matter halos. As a result, substructure lensing analyses performed with quadruple-image lenses use only the image positions and flux ratios to constrain the lens model and the properties of dark matter substructure.
@@ 2.1 The Bayesian inference problem
where $\boldsymbol{O}=\left(\boldsymbol{O_{1}},\boldsymbol{O_{2}},....\right)$ represents the data from a sample of lenses, $\pi\left({{\bf{q}}}\right)$ represents the prior on ${{\bf{q}}}$, and where $\mathcal{L}\left({{\bf{O}}}_{i}|{{\bf{q}}}\right)$ represents the likelihood function for a single lens with data ${{\bf{O}}}_{i}\equiv\left(\boldsymbol{O_{\rm{pos}}},\boldsymbol{O_{\rm{img}}},\boldsymbol{O_{\rm{f}}}\right)$. To connect the hyper-parameters ${{\bf{q}}}$ with $\boldsymbol{O_{i}}$, we must sample from an enormous volume of parameter space that specifies the masses, positions, density profiles of a typically very large $\mathcal{O}\left(10^{3}\right)$ number of dark matter subhalos and field halos. Hereafter, we will refer to ${\boldsymbol{m}_{\rm{sub}}}$ as a realization of substructure that we generate from the dark matter model specified by ${{\bf{q}}}$. The likelihood follows from marginalizing over many realizations
@@
In the preceding equation, we have introduced a vector of nuisance parameters, $\boldsymbol{\mathcal{N}}$, which include quantities that parameterize the main deflector mass profile (hereafter the macromodel), the spatial extent and structure of the lensed background quasar, and the surface brightness profile of the quasar host galaxy; $p\left(\boldsymbol{\mathcal{N}}\right)$ represents a prior on these parameters. We have included importance weights $w\left({{\bf{q}}}|\boldsymbol{O_{{i}}}\right)$, a topic we will discuss in Section 2.2.3.
The first term in the integrand, $p\left({{\bf{O}}}_{i}|{\boldsymbol{m}_{\rm{sub}}},\boldsymbol{\mathcal{N}}\right)$, is the product of an astrometric term computed with the lensed image positions $p\left(\boldsymbol{O}_{\rm{pos}}|{\boldsymbol{m}_{\rm{sub}}},\boldsymbol{\mathcal{N}}\right)$, a term that depends on the imaging data, $p\left(\boldsymbol{O}_{\rm{img}}|{\boldsymbol{m}_{\rm{sub}}},\boldsymbol{\mathcal{N}}\right)$, and the flux ratios $p\left(\boldsymbol{O}_{\rm{fr}}|{\boldsymbol{m}_{\rm{sub}}},\boldsymbol{\mathcal{N}}\right)$. We perform this integral through Monte-Carlo sampling of the parameter space, and define a likelihood through an Approximate Bayesian Computing approach, as discussed in the next section.
@@ 2.2.2 The flux ratio likelihood
In our analysis, we generate $\sim 10^{6}$ realizations per lens, and accept the top $N_{\rm{accept}}=3,000$ samples corresponding to the lowest values of $S$. This results in values of $\epsilon$ that range between 0.01 and 0.1 among the (mock) lenses in our sample. As in previous work, we handle observational measurement uncertainties by adding them post-processing to the model-predicted flux ratios before computing $S$. The convergence tests performed for ABC rejection algorithms applied to image flux ratios by Gilman et al. (2020) motivate our choice of the tolerance threshold on $\epsilon$. Finally, we obtain a continuous approximation of the flux-ratio sampling distribution marginalized over $\boldsymbol{\mathcal{N}}$, $p\left(\boldsymbol{O_{\rm{f}}}|\bf{q}_{\rm{sub}}\right)$, by applying a Gaussian kernel density estimator to the accepted samples drawn from the prior on ${{\bf{q}}}$.
@@ 2.2.3 The imaging data likelihood
We account for the imaging data constraints by applying an ABC rejection cut to a second summary statistic, $L$, which we define as the probability of observing the imaging data $L\equiv p\left(\boldsymbol{O_{\rm{img}}}|{\boldsymbol{m}_{\rm{sub}}},\boldsymbol{\mathcal{N}}\right)$. In principle, we could incorporate information directly through importance sampling with weights equal to $L$, but this causes numerical stability issues when the relative likelihood among the accepted samples fluctuate by large factors of $\sim e^{5}$ due to the high dimension of the data vector. We define the acceptance threshold on $L$ as the value of $L$ that corresponds to the top $2\%$ of the imaging data likelihoods. This choice represents a compromise between selecting models that best fit the imaging data, and accepting realizations that minimize the $S$ statistic computed with respect to the flux ratios. The optimum choice for this threshold should depend on the constraining power of the imaging data, which in turn depends on the specifics of the lens system of in question and in particular the brightness of the lensed arc.
As we show in Section 5, the imaging data enables stronger conclusions regarding substructure properties by imposing tighter constraints on the mass profile of the main deflector than one obtains from analyzing only the image positions and flux ratios. However, our method of incorporating imaging data exposes the likelihood function to a systematic bias associated with reconstruction of the source light profile with substructure in the lens model. This bias manifests as a systematic preference for lens models with less substructure. Moreover, in rare cases a dark matter halo imparts a strong perturbation to the surface brightness of a lensed arc, a feature of the data that the inference methodology we present in this work is not intended to properly capture in the likelihood. For reasons we elaborate on further in Appendix A, we conclude that the imaging data likelihood will not give a reliable inference of substructure properties on its own when calculated according to the methodology we present in this work.
To mitigate the effects of systematics associated with the imaging data in the presence of substructure, we demand that our posterior distribution have the property
@@
That is, the posterior should equal the prior on ${{\bf{q}}}$ when computed only using the imaging data (as $\epsilon\rightarrow\infty$ the flux ratio ratio likelihood term becomes uninformative). We can arrange that our posterior meets this requirement by an appropriate definition of the importance sampling weights in Equation 2. We define the importance weights, $w\left({{\bf{q}}}|\boldsymbol{O_{i}}\right)$, as
@@
$=$ $\int p\left(\boldsymbol{O_{\rm{img}}},\boldsymbol{O_{\rm{pos}}}|{\boldsymbol{m}_{\rm{sub}}},\boldsymbol{\mathcal{N}}\right)$
×$\times p\left({\boldsymbol{m}_{\rm{sub}}}|{{\bf{q}}}\right)p\left(\boldsymbol{\mathcal{N}}\right)d\boldsymbol{\mathcal{N}}d{\boldsymbol{m}_{\rm{sub}}}.$
In other words, we divide the likelihood computed for all available data, ${{\bf{O}}}_{i}$, by the imaging data and astrometric likelihood, ignoring flux ratio information. As one can easily verify, applying the importance weights in Equation 2.2.3 and performing a dark matter inference without incorporating constraints from the flux ratios yields a posterior distribution equal to $\pi\left({{\bf{q}}}\right)$.
One can interpret this term as a prior on the dark matter model ${{\bf{q}}}$ applied to each lens that depends on the parameterization of the source light profile in $\boldsymbol{\mathcal{N}}$. The source reconstruction process involves solving for a configuration of light in the source plane that maximizes the likelihood of the lensed image in the image plane. As discussed in Section 4.3, we model the source light profile as a parametric component plus a shapelet basis expansion, and we optimize the coefficients of the shapelet sets (Birrer et al., 2015) simultaneously with the parametric components. Properties of substructure can, to some degree, be absorbed by the source light profile, and the degree to which this occurs depends on various choices regarding the source light model and the process of reconstructing the extended light in the source plane (e.g. Ballard et al., 2024). The importance weights in Equation 2.2.3 reflect our knowledge that choices regarding the source light reconstruction will affect our beliefs regarding the probability of a dark matter model $\boldsymbol{{{\bf{q}}}}$. The importance weights $w\left({{\bf{q}}}|\boldsymbol{O_{\rm{img}}},\boldsymbol{O_{\rm{pos}}}\right)$ quantify this effect, and their incorporation in the likelihood corrects for the resulting bias.